1. Eaned $16,200 of cash revenue. 2. Borrowed $12,000 cash from the bank. 3. Adjusted the accounting records to recognize accrued interest expense on the bank note. The note, Issued on September 1, 2018 had a one-year term and an 8 percent annual interest rate. Required a. What is the amount of b. What amount of cash was paid for interest in 2018? c. Use a horizontal statements model to show how each event affects the balance sheet, income statement, and statement of cash expense to record for 20187 flows. Indicate whether the event increases (0 decreases (D), ar does not affect (NA) each element of the financial statements. In Cash Flows column, designate the cash flows as operating activities (OA), investing activities (IA), or financing activities (FA The transaction has been recorded as an example. Complete this question by entering your answers in the tabs below Req A and Req C Use a horizontal statements model to show how each event affects the balance sheet, income statement, and statement of cash flows decreases (D), or does not affect (NA) each element of the fnancial statements. In the Cash Flows column, designate the cash flows a activities (IA), or financing activibies (FA). The frst transaction has been recorded as an example DARBY COMPANY Horizontal Statements Model for 2018 1of4 İİİ Next
media%2Fca0%2Fca08213a-ae9d-499a-a62f-64
Answer #1

Solution a:

Amount of interest expense to record for 2018 = $12,000 * 8% *4/12 = $320

Solution b:

Amount of cash was paid for interest in 2018 = 0

Solution c:

Darby Company
Horizontal Statement Model
Event Balance Sheet Income Statement Statement of Cash Flows
Assets = Liabilities + Stockholder's Equity Revenue - Expense = Net Income
Cash = Notes Payable + Interest Payable + Common Stock + Retained Earnings
1 I = NA + NA + NA + I I - NA = I I OA
2 I = I + NA + NA + NA NA - NA = NA I FA
3 NA = NA + I + NA + D NA - I = D NA NA